OpenStudy (lilmike234):

If your starting salary were 50,000 and you recieved a 4% increase at the end of every year for 15 years, what would be the total amount in dollars, you would have earned over the first 16 years that you would have worked?

OpenStudy (lilmike234):

So my answer would be 11,250?

OpenStudy (lilmike234):

*11,250,000

OpenStudy (mathstudent55):

\(~~~a_1~~~~ +~~~~ a_2~~~+ ~~~~a_3~~~ +~ ... +~~~~~ a_{16} \) \(50,000 + 52,000 + 54,080 +~ ...+ 90,047.17\) \(\Large S = a_1\times \dfrac{1 - r^n}{1 - r} \) where \(a_1\) = the first term r = the common ratio, and n = the number of terms \(\Large S = 50,000 \times \dfrac{1 - (1.04)^{16}}{1 - 1.04}\) \(\Large S = 1,091,226.56\)